# FedALT: Federated Fine-Tuning through Adaptive Local Training with Rest-of-World LoRA

### v2

@@ 1 Introduction
Figure 1: Illustration of FedALT. Instead of directly aggregating local LoRA modules from each client using FedAvg, FedALT introduces a frozen RoW LoRA component to transmit shared global knowledge while preserving client-specific adaptations through Individual LoRA. The adaptive mixer dynamically combines the RoW LoRA and Individual LoRA. See details in section 3.
@@
Large language models (LLMs) [19, 4, 31, 37, 50, 48, 38] have demonstrated exceptional capabilities in language understanding and generation, enabling a wide range of applications. Adapting pretrained LLMs to specialized domains or further enhancing their performance requires fine-tuning with task-specific data [29]. However, full fine-tuning, which involves updating all model parameters, is computationally expensive and often impractical for real-world deployment. To address this limitation, several parameter-efficient fine-tuning (PEFT) [9, 11, 14, 25, 24, 23] methods have been proposed, with low-rank adaptation (LoRA) [16] being one of the most widely used. LoRA reduces the number of trainable parameters by integrating low-rank matrices into the model, significantly reducing computational costs.
@@
Fine-tuning LLMs necessitates substantial data volumes for adaptation to downstream tasks; however, this requisite data often resides across multiple institutions where privacy regulations and security protocols prohibit direct data sharing. Federated Learning (FL) [27] facilitates collaborative fine-tuning without raw data exposure, establishing itself as an essential paradigm for privacy-preserving adaptation. Nonetheless, existing federated LoRA fine-tuning methodologies [49, 34, 43, 12, 3] predominantly presuppose a singular global model, neglecting the inherent data heterogeneity among participating clients. Such a homogeneous modeling approach frequently yields suboptimal performance and introduces fairness disparities, particularly in real-world deployments where client data exhibits significant variance in volume, domain, and distributional characteristics. While limited research on personalized federated LoRA fine-tuning [15, 30] exists, these approaches primarily address label heterogeneity cases. However, in realistic LLM fine-tuning contexts, clients’ learning objectives naturally diverge due to diverse business imperatives [47]. Personalized federated LoRA fine-tuning addressing more realistic task-heterogeneous scenarios [45] remains insufficiently investigated in the current literature.
@@
Federated Averaging (FedAvg) [27] is perhaps the most widely used FL framework, serving as the foundation for numerous follow-up works that introduce various optimizations. Its core principle is to aggregate locally trained client models into a global model, which then initializes the next round of local training. Most personalized FL algorithms [45, 30, 35, 8, 7, 10], despite their implementation differences, build upon the FedAvg framework by incorporating personalization mechanisms—and personalized federated LoRA fine-tuning follows the same paradigm. For example, the recently proposed FedDPA [45] algorithm personalizes LoRA by introducing both global and local LoRA components. The global LoRA component is trained using FedAvg, while the local LoRA component is trained independently on each client’s private data. This local adaptation can occur either after the global LoRA stabilizes or alternatedly with global LoRA training, in which case the final output is computed as a weighted combination of both components using a pre-defined weighting factor. However, our motivational study (see section 2.3) reveals that FedAvg-based approaches may degrade client-specific performance due to harmful cross-client interference. We identify two fundamental limitations of FedAvg-based methods like FedDPA in heterogeneous FL settings. First, they suffer from harmful cross-client interference due to reliance on global aggregation, which can negate individual client improvements achieved during local fine-tuning. Second, these methods lack an effective mechanism for balancing global and local information, relying on fixed or predefined weighting factors, which can be suboptimal for clients poorly represented by the global model.
@@
To address these critical limitations, we propose a novel personalized federated LoRA fine-tuning approach called FedALT (short for Federated Fine-tuning through Adaptive Local Training with Rest-of-World LoRA.). Distinctly different from FedAvg-based methods, FedALT does not aggregate local models to initialize training rounds. Instead, each client continues training on their own previously trained local models, supplemented by a frozen Rest-of-World (RoW) LoRA component. The RoW LoRA captures global knowledge without updating during local training rounds, thus avoiding interference with client-specific adaptations. To leverage RoW LoRA effectively, we introduce an adaptive mixer, inspired by Mixture-of-Experts (MoE) [18], which dynamically and optimally combines global and local information for each client’s data sample. Our contributions are as follows:
@@
  • We perform comprehensive experiments on two widely-used LLMs (Bloom [21] and Llama 2 [38]) using the Flan benchmark [6], demonstrating that FedALT significantly outperforms existing federated LoRA fine-tuning methods.
2 Preliminaries and Motivations
2.1 LoRA
@@
where $\mathbf{A}$ is typically initialized with random Gaussian values, while $\mathbf{B}$ is initialized to zero, ensuring a stable start to the fine-tuning process.
2.2 Heterogeneous Federated Fine-Tuning
@@
In this scenario, each client $k$ aims to leverage FL to extract useful information while fine-tuning a model tailored to its specific task. The goal is for each client to learn an individualized fine-tuned model $\textbf{W}_{k}$ that best fits its local dataset. The objective of the training process can be formulated as:
@@
where $L_{k}(\cdot)$ represents the loss function for client $k$, and $\mathcal{W}$ denotes the set of personalized fine-tuned models $\{\mathbf{W}_{k}\}_{k=1}^{K}$.
2.3 Motivational Study
@@
In FL settings, the data collected by each client can be highly diverse [17, 28, 40]. We conducted a motivational experiment using the Flan dataset [6], selecting 8 tasks and creating 8 clients, each specializing in a single task. To demonstrate potential interference in federated fine-tuning, we compared two baseline methods. The first method, FedIT [49], combines FedAvg with LoRA fine-tuning. The second method is Local Only LoRA fine-tuning, where each client fine-tunes its LoRA locally without any communication or information sharing between clients.
Figure 2: Motivational study results. See details in section 2.3.
The results, shown in fig. 2, support our first observation. On one hand, federated fine-tuning using FedIT achieves better performance than local fine-tuning on tasks Coreference Resolution and Paraphrase Detection, confirming that FL can enhance learning by enabling clients to benefit from shared knowledge. On the other hand, for tasks Commonsense Reasoning and Text Classification, FedIT underperforms compared to individual local fine-tuning. This dual outcome underscores both the promise and the pitfalls of federated fine-tuning in heterogeneous settings.
@@
Having established the existence of harmful interference, we next explored potential mitigation strategies. In centralized LoRA fine-tuning, HydraLoRA [36] found that splitting a single large LoRA into multiple smaller LoRAs reduces cross-domain interference. However, FL introduces additional challenges, as data domains are distributed across clients. To test whether multiple smaller LoRAs could mitigate interference in FL, we implemented FedAvg with multiple parallel LoRAs (i.e. FedIT-split). However, as shown in fig. 2, its performance remained comparable to FedIT (FedAvg with a single large LoRA), indicating that interference persisted. This result is expected: in centralized settings, interference occurs within a shared dataset and can be mitigated by partitioning LoRAs. In FL, interference arises primarily during server aggregation, and simply introducing multiple LoRAs does not prevent the aggregated model from disrupting local adaptations. Consequently, direct adaptations of centralized interference-mitigation strategies fail to address the core issue in FL.
3 Proposed Method: FedALT
@@
Our motivational studies reveal that harmful interference in FL arises from server aggregation and the use of the aggregated model to initialize subsequent training rounds. In the FedAvg paradigm, a global LoRA serves as a shared initialization point, allowing clients to fine-tune locally on their respective datasets. After fine-tuning, the updated LoRAs are averaged on the server and redistributed to clients as the new initialization. However, when client data differ significantly, aggregation disrupts individual adaptations, often negating improvements from local training and hindering personalization. Can cross-client interference be mitigated? The simplest and most direct solution is local fine-tuning using only a client’s own dataset. This approach eliminates interference entirely but fails to leverage shared knowledge from other clients. To address this trade-off, we propose a novel personalized federated LoRA fine-tuning algorithm that fundamentally diverges from the FedAvg paradigm, prioritizing local training while selectively integrating useful global information.
@@
Our method introduces two key innovations: 1. Each client maintains two LoRA modules: an Individual LoRA, which captures locally learned information and is updated exclusively during the client’s fine-tuning, and a Rest-of-World (RoW) LoRA, which aggregates information from all other clients. Crucially, the RoW LoRA remains fixed during local fine-tuning, ensuring that aggregation does not interfere with a client’s adaptation. 2. An adaptive mixer dynamically learns the optimal input-specific weighting between the Individual LoRA and RoW LoRA components to balance personalization and global knowledge integration. In the following of this section, we detail the design and functionality of these components. Additionally, a theoretical convergence analysis of the proposed method is provided in the appendix F.
3.1 Individual and Rest-of-World (RoW) LoRA
@@
Each client $k$ maintains two types of LoRAs: an Individual LoRA, denoted by $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$, and a Rest-of-World (RoW) LoRA, denoted by $\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$. The Individual LoRA captures information obtained locally by the client and is isolated from the information of other clients. In contrast, the RoW LoRA represents the average of the Individual LoRAs from all other clients and remains frozen during the client’s local fine-tuning. Specifically, at the end of each round, client $k$ updates its Individual LoRA $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$ through local training. The RoW LoRA for client $k$ is then computed as:
@@
Now, consider the scenario where a weight $\frac{1}{K}$ is assigned to the local LoRA and $1-\frac{1}{K}$ to the RoW LoRA when combining their outputs. In this case, the combined LoRA is mathematically equivalent to the averaged LoRA. This equivalence arises naturally from the definitions of the RoW and averaged LoRAs. If these two LoRAs are trained simultaneously on local datasets, the approach essentially reduces to the standard FedAvg (or FedIT) algorithm, as all clients would end up with the same combined LoRA.
However, it is crucial to emphasize that the RoW LoRA is not further updated on clients using their local datasets. Instead, it is only updated through averaging at the end of each learning round. Training the RoW LoRA on local datasets is unnecessary and potentially counterproductive, as subsequent averaging could cancel out any improvements. Furthermore, skipping local training for the RoW LoRA reduces the computational workload on clients by half.
3.2 Dynamic Mixture-of-Experts Mechanism
@@
Given the complexity of modern LLMs, where LoRA modules are inserted at multiple layers, learning $\alpha(x)$ is non-trivial. Inspired by [18], we adopt a Mixture-of-Experts (MoE) approach to compute $\alpha(x)$. Specifically, we introduce a mixer $\mathbf{G}_{k}\in\mathbb{R}^{2\times d}$ which is a dense layer with trainable weights (a transformation matrix), followed by a softmax function, to learn the weighting:
@@
3.3 Algorithm Workflow
By combining the two LoRA modules (Individual LoRA and RoW LoRA) with dynamic weights, our proposed method achieves a balance between leveraging useful information from other clients and isolating harmful interference. To better illustrate the FedALT training process, we provide an overview of the server and client operations during fine-tuning at round $t$.
@@ 3.3 Algorithm Workflow
The theoretical analysis of the convergence of the proposed FedALT is presented in appendix F.
3.4 Discussion: Why a Single RoW LoRA?
4 Experiments
In this section, we evaluate and compare the performance of our proposed method, FedALT, with other baseline methods. Detailed training implementations are provided in the appendix B.
Datasets. Our training datasets are derived from Flan [6], a comprehensive benchmark comprising over 60 NLP datasets across more than 12 diverse task types. To simulate client heterogeneity in the main experiments, we assume 8 clients, each assigned a distinct NLP task dataset. Sensitivity analysis with varying numbers of clients and degrees of task heterogeneity are presented in section 4.3. Detailed dataset descriptions, including task distributions and client assignments, are provided in appendix A.
Pre-trained Model and Baselines. We utilize LLaMA2-7B [38] and Bloom-560M [21] as the pre-trained models for all baseline and proposed fine-tuning methods. To evaluate the performance of our proposed method, we compare it with the state-of-the-art methods in federated fine-tuning with LoRA, including FedIT [49], FFA-LoRA [34]. Additionally, we compare it with state-of-the-art methods focused on personalized federated fine-tuning with LoRA, such as FedDPA [45], PF2LoRA [15], FedSA [12], and FDLoRA [30]. We also include the Local Only baseline, where clients fine-tune their models independently without federated collaboration. Detailed descriptions of these baseline methods are provided in the appendix C.
Table 1: Performance comparison with baseline methods with LLaMA2-7B. The average column represents the overall performance across these tasks. See details in section 4.1.
Methods LLaMA2-7B
Commonsense Coreference Natural Language Paraphrase Reading Sentiment Struct to Text Average
@@ 4 Experiments
FedALT 76.12 83.04 67.73 77.80 59.41 51.57 53.12 71.60 67.55
Table 2: Performance comparison across various NLP tasks with Bloom-560M. The average column represents the overall performance across these tasks. See details in section 4.1.
Methods Bloom-560M
Commonsense Coreference Natural Language Paraphrase Reading Sentiment Struct to Text Average
@@
FedALT 56.39 41.27 44.45 70.67 30.62 43.04 39.24 59.03 48.09
4.1 Performance Comparisons
@@
4.2 Ablation Study
Table 3: Impact of decoupling LoRA training. See Sec. 4.2.
@@ 4.2 Ablation Study
Impact of Decoupling LoRA Training. In our proposed method, only the Individual LoRA modules are updated locally during client fine-tuning, while the RoW LoRA remains frozen and is updated exclusively through server-side aggregation. This design mitigates harmful cross-client interference by preventing local updates to the shared global component. To evaluate the effectiveness of this decoupled training strategy, we compare FedALT with two alternative methods. The first baseline, RoW-Update (rank = 4), fine-tunes both the Individual and RoW LoRA modules locally, each with a rank of 4 to maintain the same total number of trainable parameters. The second baseline, RoW-Update (rank = 8), also fine-tunes both LoRA modules locally, but assigns a rank of 8 to each, resulting in a higher overall parameter count. For fair comparison, both baselines use the dynamic mixer. The key distinction lies in whether the RoW LoRA is updated locally. As shown in table 3, while the second method outperforms the first due to its increased capacity, both alternatives underperform relative to FedALT. Notably, even with more trainable parameters, the second method fails to match FedALT’s performance, highlighting the benefits of decoupling local updates from the global component.
@@
Impact of Dynamic Weighting via MOE. To enable input-specific flexibility, our proposed method adopts a MOE mechanism to dynamically adjust the contributions of the Individual LoRA and RoW LoRA for each data sample. We conduct experiments to evaluate the efficiency of this dynamic weighting by comparing it to two fixed weighting strategies. In the first fixed strategy, we set the contribution weight to a fixed value of $\alpha=0.5$, meaning that the Individual LoRA and RoW LoRA contribute equally. In the second fixed strategy, we set the contribution weight to $\alpha=1/K$, where $K$ is the number of clients. This configuration effectively reduces the method to using the averaged LoRA, relying solely on each client’s last-round LoRA modules for the next round’s initialization. The results in fig. 3(b) validate that our proposed dynamic weighting mechanism outperforms both fixed strategies, demonstrating its ability to enhance performance. Additionally, our dynamic weighting is implemented through the personal $\mathbf{G}_{k}$ mixer, which is maintained locally by each client and is not aggregated during the training process. To further investigate this design choice, we conducted experiments where $\mathbf{G}_{k}$ is averaged across clients. The results confirm that retaining a personalized $\mathbf{G}_{k}$ for each client yields optimal performance, emphasizing the importance of input-specific and client-specific adjustments.
(a) Impact of RoW LoRA. Isolating client-specific information is more effective than using globally averaged updates. See section 4.2.
(b) Impact of dynamic weighting via MOE. The personalized mixer outperforms fixed or shared weighting. See section 4.2.
Figure 3: Ablation studies of FedALT.
4.3 Sensitivity Analysis
Effect of Data Heterogeneity. To further evaluate the robustness of our method under varying degrees of client heterogeneity, we design experiments across three settings. Our main experimental configuration, where each client is assigned a unique task, is denoted as the High Heterogeneous setting. We also include two alternative configurations. In the Low Heterogeneous setting, 8 clients are divided between only two tasks—Sentiment Analysis and Natural Language Inference—with clients sharing a task drawing data from the same source, resulting in minimal heterogeneity. The Mild Heterogeneous setting involves 8 clients assigned datasets spanning a total of 6 distinct NLP tasks. As shown in table 4, FedALT consistently outperforms all baseline methods across all levels of heterogeneity, using both Bloom and LLaMA2 as pre-trained models. These results highlight the robustness and adaptability of FedALT in handling diverse and heterogeneous data distributions. Due to space limitations, we report only the average results here; detailed per-client performance can be found in the appendix E.

### v3

@@ Introduction
Figure 1: Illustration of FedALT. Instead of directly aggregating local LoRA modules from each client using FedAvg, FedALT introduces a frozen RoW LoRA component to transmit shared global knowledge while preserving client-specific adaptations through Individual LoRA. The adaptive mixer dynamically combines the RoW LoRA and Individual LoRA.
@@
Large language models (LLMs) (kenton2019bert; brown2020language; raffel2020exploring; touvron2023llama; zhou2024comprehensive; zeng2022glm; touvron2023llama2) have demonstrated exceptional capabilities in language understanding and generation, enabling a wide range of applications. Adapting pretrained LLMs to specialized domains or further enhancing their performance requires fine-tuning with task-specific data (minaee2024large). However, full fine-tuning, which involves updating all model parameters, is computationally expensive and often impractical for real-world deployment. To address this limitation, several parameter-efficient fine-tuning (PEFT) (ding2023parameter; fu2023effectiveness; liu2022few; li2021prefix; lester2021power) methods have been proposed, with low-rank adaptation (LoRA) (hu2021lora) being one of the most widely used. LoRA reduces the number of trainable parameters by integrating low-rank matrices into the model, significantly reducing computational costs.
@@
Fine-tuning LLMs necessitates substantial data volumes for adaptation to downstream tasks; however, this requisite data often resides across multiple institutions where privacy regulations and security protocols prohibit direct data sharing. Federated Learning (FL) facilitates collaborative fine-tuning without raw data exposure, establishing itself as an essential paradigm for privacy-preserving adaptation (bian2025survey). Nonetheless, existing federated LoRA fine-tuning methodologies (zhang2024towards; sun2024improving; wang2024flora; Bian_2025_ICCV) predominantly presuppose a singular global model, neglecting the inherent data heterogeneity among participating clients. Such a homogeneous modeling approach frequently yields suboptimal performance and introduces fairness disparities, particularly in real-world deployments where client data exhibits significant variance in volume, domain, and distributional characteristics. While limited research on personalized federated LoRA fine-tuning (hao2025personalized; qi2024fdlora) exists, these approaches primarily address label heterogeneity cases. However, in realistic LLM fine-tuning contexts, clients’ learning objectives naturally diverge due to diverse business imperatives (yao2022benchmark). Personalized federated LoRA fine-tuning addressing more realistic task-heterogeneous scenarios (yang2024dual) remains insufficiently investigated in the current literature.
@@
Federated Averaging (FedAvg) (mcmahan2017communication) is perhaps the most widely used FL framework, serving as the foundation for numerous follow-up works that introduce various optimizations. Its core principle is to aggregate locally trained client models into a global model, which then initializes the next round of local training. Most personalized FL algorithms (yang2024dual; qi2024fdlora; tan2022towards; deng2020adaptive; collins2021exploiting; fallah2020personalized), despite their implementation differences, build upon the FedAvg framework by incorporating personalization mechanisms—and personalized federated LoRA fine-tuning follows the same paradigm. For example, the recently proposed FedDPA (yang2024dual) algorithm personalizes LoRA by introducing both global and local LoRA components. The global LoRA component is trained using FedAvg, while the local LoRA component is trained independently on each client’s private data. This local adaptation can occur either after the global LoRA stabilizes or alternatedly with global LoRA training, in which case the final output is computed as a weighted combination of both components using a pre-defined weighting factor. However, our motivational study reveals that FedAvg-based approaches may degrade client-specific performance due to harmful cross-client interference. We identify two fundamental limitations of FedAvg-based methods like FedDPA in heterogeneous FL settings. First, they suffer from harmful cross-client interference due to reliance on global aggregation, which can negate individual client improvements achieved during local fine-tuning. Second, these methods lack an effective mechanism for balancing global and local information, relying on fixed or predefined weighting factors, which can be suboptimal for clients poorly represented by the global model.
@@
To address these critical limitations, we propose a novel personalized federated LoRA fine-tuning approach called FedALT (short for Federated Fine-tuning through Adaptive Local Training with Rest-of-World LoRA.). Distinctly different from FedAvg-based methods, FedALT does not aggregate local models to initialize training rounds. Instead, each client continues training on it own previously trained local models, supplemented by a frozen Rest-of-World (RoW) LoRA component. The RoW LoRA captures global knowledge without updating during local training rounds, thus avoiding interference with client-specific adaptations. To leverage RoW LoRA effectively, we introduce an adaptive mixer, inspired by Mixture-of-Experts (MoE) (jordan1994hierarchical), which dynamically and optimally combines global and local information for each client’s data sample. Our contributions are as follows:
@@
  • We perform comprehensive experiments on two LLMs (Bloom (le2023bloom) and Llama 2 (touvron2023llama2)) using the Flan benchmark (chung2024scaling), demonstrating that FedALT significantly outperforms existing federated LoRA fine-tuning methods.
Related Work
Parameter-Efficient Fine-Tuning
As the size of LLMs continues to grow, full parameter fine-tuning has become increasingly computationally and storage-intensive (hadi2023survey). To address this issue, Parameter-Efficient Fine-Tuning (PEFT) methods (ding2023parameter; fu2023effectiveness; han2024parameter; liu2022few; li2021prefix; lester2021power) have been developed to significantly reduce the number of trainable parameters. PEFT techniques introduce a limited set of additional trainable parameters to enhance model performance while keeping the majority of pre-trained parameters frozen. Among all existing PEFT methods, the most prominent approach is LoRA (hu2021lora), which employs low-rank matrices to approximate the pre-trained weight matrix, updating only the low-rank components. LoRA has became a standard method for adapting LLMs like Llama under limited computational resources (lermen2023lora). Several works have been proposed to improve LoRA (tian2024hydralora; lermen2023lora; wang2023multilora; sheng2023s; liu2023moelora). A recent development, Hydra-LoRA (tian2024hydralora), employs an asymmetric structure with a shared matrix for all samples and distinct matrices for each intrinsic component, thereby enhancing domain adaptation. While these advancements primarily focus on centralized learning settings, our work extends the application of PEFT methods to federated LLM fine-tuning, addressing the unique challenges posed by distributed and heterogeneous client data.
Federated Learning
The foundational work, FedAvg (mcmahan2017communication), addresses privacy and communication efficiency by aggregating local model to train a shared global model. Since then, numerous studies (liuimproving; wang2024aggregation; yang2024fedfed; liu2025fedadamwcommunicationefficientoptimizerconvergence; bian2024accelerating; zhang2025fedel; 10.1145/3746027.3755226; liu2024fedbcgd) have focused on tackling various challenges within FL settings. One of the key challenges is data heterogeneity, which makes it difficult to train a single shared global model that performs well across all clients. To address this, Personalized Federated Learning (PFL) (tan2022towards; deng2020adaptive; collins2021exploiting; fallah2020personalized) has emerged as an approach to adapt the global model to the specific needs of each client. Most PFL research has focused on addressing statistical differences in data distributions (e.g., label distribution skew) and has been applied primarily to smaller, standard models. In contrast, our work addresses the unique challenges posed by heterogeneous clients with diverse tasks by leveraging LLM fine-tuning within the FL framework.
Federated Fine-Tuning
Several studies (cho2024heterogeneous; wang2024flora; kuang2024federatedscope; wu2024fedbiot) have explored federated fine-tuning approaches. (kuang2024federatedscope) proposed federated full parameters fine-tuning, while (sun2022conquering) introduced federated fine-tuning with PEFT using prefix-tuning. (zhang2024towards) was the first study to apply LoRA in a federated context. (bai2024federated; cho2024heterogeneous) focused on federated fine-tuning where clients have different LoRA ranks, while (wang2024flora) proposed strategies for LoRA initialization to achieve better performance. Another studies (Bian_2025_ICCV; sun2024improving) addressed server aggregation bias in LoRA-based federated fine-tuning.
The most relevant works to ours focus on personalized federated fine-tuning with LoRA (yang2024dual; qi2024fdlora). FedDPA (yang2024dual) introduced global and local LoRA components but sufferd from interference during global LoRA training and struggled to balance the two phases effectively. PF2LoRA (hao2025personalized) employed a bilevel framework to combine a shared adapter with client-specific adapters but faced interference issues during the aggregation of the shared adapter. FDLoRA (qi2024fdlora) utilized personalized LoRA modules to initialize federated training, combining them with a global LoRA module via adaptive fusion; however, it relied on server-side datasets and was prone to interference during global LoRA aggregation, limiting its effectiveness in heterogeneous settings.
Preliminaries and Motivations
Low-Rank Adaptation (LoRA)
@@
where $\mathbf{A}$ is typically initialized with random Gaussian values, while $\mathbf{B}$ is initialized to zero.
Heterogeneous Federated Fine-Tuning
@@
In this scenario, each client $k$ aims to leverage FL to extract useful information while fine-tuning a model tailored to its specific task. The goal is for each client to learn an individualized fine-tuned model $\textbf{W}_{k}$ that best fits its local dataset. The objective can be formulated as:
@@
where $L_{k}(\cdot)$ represents the loss function for client $k$, and $\mathcal{W}$ denotes the set of personalized models $\{\mathbf{W}_{k}\}_{k=1}^{K}$.
Motivational Study
@@
In FL settings, the data collected by each client can be highly diverse (huang2022learn; mendieta2022local; wang2024taming). We conducted a motivational experiment using the Flan dataset (chung2024scaling), selecting 8 tasks and creating 8 clients, each specializing in a single task. To demonstrate potential interference in federated fine-tuning, we compared two baseline methods. The first method, FedIT (zhang2024towards), combines FedAvg with LoRA fine-tuning. The second method is Local Only LoRA fine-tuning, where each client fine-tunes its LoRA locally without any communication or information sharing between clients. The results, shown in fig. 2, support our first observation. On one hand, federated fine-tuning using FedIT achieves better performance than local fine-tuning on tasks Coreference Resolution and Paraphrase Detection, confirming that FL can enhance learning by enabling clients to benefit from shared knowledge. On the other hand, for tasks Commonsense Reasoning and Text Classification, FedIT underperforms compared to individual local fine-tuning. This dual outcome underscores both the promise and the pitfalls of federated fine-tuning in heterogeneous settings.
Figure 2: Motivational study results.
@@
Having established the existence of harmful interference, we next explored potential mitigation strategies. In centralized LoRA fine-tuning, HydraLoRA (tian2024hydralora) found that splitting a single large LoRA into multiple smaller LoRAs reduces cross-domain interference. However, FL introduces additional challenges, as data domains are distributed across clients. To test whether multiple smaller LoRAs could mitigate interference in FL, we implemented FedAvg with multiple parallel LoRAs (i.e. FedIT-split). However, as shown in fig. 2, its performance remained comparable to FedIT (FedAvg with a single large LoRA), indicating that interference persisted. This result is expected: in centralized settings, interference occurs within a shared dataset and can be mitigated by partitioning LoRAs. In FL, interference arises primarily during server aggregation, and simply introducing multiple LoRAs does not prevent the aggregated model from disrupting local adaptations. Consequently, direct adaptations of centralized interference-mitigation strategies fail to address the core issue in FL.
Proposed Method: FedALT
@@
Our motivational studies reveal that harmful interference in FL arises from server aggregation and the use of the aggregated model to initialize subsequent training rounds. In the FedAvg paradigm, a global LoRA serves as a shared initialization point, allowing clients to fine-tune locally on their respective datasets. After fine-tuning, the updated LoRAs are averaged on the server and redistributed to clients as the new initialization. However, when client data differ significantly, aggregation disrupts individual adaptations, often negating improvements from local training and hindering personalization. Can cross-client interference be mitigated? The simplest and most direct solution is local fine-tuning using only a client’s own dataset. This approach eliminates interference entirely but fails to leverage shared knowledge from other clients. To address this trade-off, we propose a novel personalized federated LoRA fine-tuning algorithm that diverges from the FedAvg paradigm, prioritizing local training while selectively integrating useful global information.
@@
Our method introduces two key innovations: 1. Each client maintains two LoRA modules: an Individual LoRA, which captures locally learned information and is updated exclusively during the client’s fine-tuning, and a Rest-of-World (RoW) LoRA, which aggregates information from all other clients. Crucially, the RoW LoRA remains fixed during local fine-tuning, ensuring that aggregation does not interfere with a client’s adaptation. 2. An adaptive mixer dynamically learns the optimal input-specific weighting between the Individual LoRA and RoW LoRA components to balance personalization and global knowledge integration. In the following of this section, we detail the design and functionality of these components. The theoretical convergence analysis is provided in the supplementary material.
Individual and Rest-of-World (RoW) LoRA
@@
Each client $k$ maintains two types of LoRAs: an Individual LoRA, denoted by $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$, and a RoW LoRA, denoted by $\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$. The Individual LoRA captures information obtained locally by the client and is isolated from the information of other clients. In contrast, the RoW LoRA represents the average of the Individual LoRAs from all other clients and remains frozen during the client’s local fine-tuning. Specifically, at the end of each round, client $k$ updates its Individual LoRA $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$ through local training. The RoW LoRA for client $k$ is then computed as:
@@
It is crucial to emphasize that the RoW LoRA is not further updated on clients using their local datasets. Instead, it is only updated through averaging at the end of each learning round. Training the RoW LoRA on local datasets is unnecessary and potentially counterproductive, as subsequent averaging could cancel out any improvements. Furthermore, skipping local training for the RoW LoRA reduces the computational workload on clients by half.
Dynamic Mixture-of-Experts Mechanism
@@
Given the complexity of modern LLMs, where LoRA modules are inserted at multiple layers, learning $\alpha(x)$ is non-trivial. Inspired by (jordan1994hierarchical), we adopt a Mixture-of-Experts (MoE) approach to compute $\alpha(x)$. Specifically, we introduce a mixer $\mathbf{G}_{k}\in\mathbb{R}^{2\times d}$ which is a dense layer with trainable weights (a transformation matrix), followed by a softmax function, to learn the weighting:
@@
Algorithm Workflow
By combining the two LoRA modules (Individual LoRA and RoW LoRA) with dynamic weights, our proposed method achieves a balance between leveraging useful information from other clients and isolating harmful interference. To better illustrate the FedALT training process, we provide an overview of the server and client operations at round $t$.
@@
Methods Commonsense Coreference Natural Language Paraphrase Reading Sentiment Struct to Text Average
@@
[…]
Bloom-560M
@@
[…]
Table 1: Performance comparison with baseline methods across different models.
Experiments
Datasets. Our training datasets are derived from Flan (chung2024scaling), a comprehensive benchmark comprising over 60 NLP datasets across more than 12 diverse task types. To simulate client heterogeneity in the main experiments, we assume 8 clients, each assigned a distinct NLP task dataset. Detailed descriptions, including task distributions and client assignments, are provided in supplementary material.
Pre-trained Model and Baselines. We utilize LLaMA2-7B (touvron2023llama2) and Bloom-560M (le2023bloom) as the pre-trained models for all baseline and proposed fine-tuning methods. To evaluate the performance of our proposed method, we compare it with the state-of-the-art methods in federated fine-tuning with LoRA, including FedIT (zhang2024towards), FFA-LoRA (sun2024improving). Additionally, we compare it with state-of-the-art methods focused on personalized federated fine-tuning with LoRA, such as FedDPA (yang2024dual), PF2LoRA (hao2025personalized), FedSA (guo2024selective), and FDLoRA (qi2024fdlora). We also include the Local Only baseline, where clients fine-tune their models independently without federated collaboration. Detailed descriptions of these baseline methods are provided in the supplementary material.
(a) Impact of RoW LoRA. Isolating client-specific information is more effective than using globally averaged updates.
(b) Impact of dynamic weighting via MOE. The personalized mixer outperforms fixed or shared weighting.
Figure 3: Ablation studies of FedALT.
(a) Effect of number of clients.
(b) Effect of LoRA rank.
(c) Effect of local epochs.
Performance Comparisons
@@
Ablation Study
@@
Table 2: Impact of decoupling LoRA training.
Impact of Decoupling LoRA Training. In our proposed method, only the Individual LoRA modules are updated locally during client fine-tuning, while the RoW LoRA remains frozen and is updated exclusively through server-side aggregation. This design mitigates harmful cross-client interference by preventing local updates to the shared global component. To evaluate the effectiveness of this decoupled training strategy, we compare FedALT with two alternative methods. The first baseline, RoW-Update (rank = 4), fine-tunes both the Individual and RoW LoRA modules locally, each with a rank of 4 to maintain the same total number of trainable parameters. The second baseline, RoW-Update (rank = 8), also fine-tunes both LoRA modules locally, but assigns a rank of 8 to each, resulting in a higher overall parameter count. For fair comparison, both baselines use the dynamic mixer. The key distinction lies in whether the RoW LoRA is updated locally. As shown in table 2, while the second method outperforms the first due to its increased capacity, both alternatives underperform to FedALT. Notably, even with more trainable parameters, the second method fails to match FedALT’s performance, highlighting the benefits of decoupling local updates from the global component.
@@
Impact of Dynamic Weighting via MOE. To enable input-specific flexibility, our proposed method adopts a MOE mechanism to dynamically adjust the contributions of the Individual LoRA and RoW LoRA for each data sample. We conduct experiments to evaluate the efficiency of this dynamic weighting by comparing it to two fixed weighting strategies. In the first fixed strategy, we set the contribution weight to a fixed value of $\alpha=0.5$, meaning that the Individual LoRA and RoW LoRA contribute equally. In the second fixed strategy, we set the contribution weight to $\alpha=1/K$, where $K$ is the number of clients. The results in fig. 3(b) validate that our proposed dynamic weighting mechanism outperforms both fixed strategies, demonstrating its ability to enhance performance. Additionally, our dynamic weighting is implemented through the personal $\mathbf{G}_{k}$ mixer, which is maintained locally by each client and is not aggregated during the training process. To further investigate this design choice, we conducted experiments where $\mathbf{G}_{k}$ is averaged across clients. The results confirm that retaining a personalized $\mathbf{G}_{k}$ for each client yields optimal performance, emphasizing the importance of input-specific and client-specific adjustments.
Sensitivity Analysis